OpenStudy (anonymous):

I suck at these don't give me the answer just help me threw it please thanks :)

OpenStudy (anonymous):



OpenStudy (anonymous):

(x-2)=4x for x

OpenStudy (johnweldon1993):

Are you stating that 2/3 is your answer? because you are almost correct x - 2 = 4x ...solve for x we subtract x from both sides x - 2 = 4x -x -x -2 = 3x and to isolate 'x'...we find that the answer is...?

OpenStudy (anonymous):

is it 2/5

OpenStudy (johnweldon1993):

not quite.... When we are down to -2 = 3x we want to isolate 'x'...and we do this by dividing everything by 3... because 3x / 3 = x right? so when we do that -2 = 3x -------- 3 3 we find that -2/3 = x x = -2/3

OpenStudy (anonymous):

yea ok

OpenStudy (johnweldon1993):

Hope that was understandable ?

OpenStudy (anonymous):

not really didnt help me with my question but is okay john lol

OpenStudy (anonymous):

these were some of the multiple chooses -2/5 -5/2 2/5 and 5

jimthompson5910 (jim_thompson5910):

if -2/3 isn't an answer choice, then there has to be a typo somewhere

OpenStudy (anonymous):

nvm jim i got it ... ;/

jimthompson5910 (jim_thompson5910):

ok great, glad you did